OpenStudy (anonymous):

If a line segment has endpoints A(3x + 5, 3y) and B(x − 1, −y), what are the coordinates of the midpoint of AB ? (1 point) (x + 3, 2y) (2x + 3, y) (2x + 2, y) (4x + 4, 2y)

OpenStudy (mertsj):

Add the x values and divide by 2 then add the y values and divide by 2

OpenStudy (mertsj):

(3x+5+x-1)/2=(4x+4)/2=2x+2 (3y+-y)/2=(2y)/2=y

OpenStudy (anonymous):

is the answer (2x+2,y)